OpenStudy (anonymous):

solve the inequality. 8| x + 3 over 4| <2

OpenStudy (candy13106):

ok um well

OpenStudy (anonymous):

im waiting

OpenStudy (candy13106):

wat does the mark in front of the 2 mean

OpenStudy (candy13106):

oh gosh heres thta guy again in ht ewhite shirt

OpenStudy (debbieg):

Isolate the abs value expression first.

OpenStudy (candy13106):

ok well wat does that slash mark or watever that is means

OpenStudy (debbieg):

You need something of the form: |something|<something

OpenStudy (debbieg):

\[\Large 8\left| \frac{ x+3 }{ 4 } \right|<2\] Divide both sides by 8: \[\Large \left| \frac{ x+3 }{ 4 } \right|<\dfrac{1}{4}\]

OpenStudy (candy13106):

ummmmmmmmmmmmmm wat does that funky thing mean plz tell come on man

OpenStudy (anonymous):

whats the answer debbie

OpenStudy (debbieg):

Really? the less than sign??

OpenStudy (candy13106):

well i know it

OpenStudy (anonymous):

what is it

OpenStudy (debbieg):

I won't give you the answer. I will help you learn how to do the problem, if you would like. If not, I'll move on. up to you.

OpenStudy (candy13106):

okok i got it i really did i know wat that means

OpenStudy (anonymous):

candy help me out, debbie ok go away then

OpenStudy (debbieg):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution. Have homework problems? We can help you - as long as you are looking to learn how to get your solution (and not just the answer). http://openstudy.com/code-of-conduct

OpenStudy (debbieg):

ok, good luck on the test.

OpenStudy (candy13106):

ok gowingspans the firt thing u do is that ......................

OpenStudy (anonymous):

I won't use this stupid stuff after college and getting my degree bye bye

OpenStudy (candy13106):

ok x= to + 3over 4 right is that wat u said in the question and have u realised that im way younger than u by a lot

OpenStudy (candy13106):

ok u there

OpenStudy (candy13106):

im waiting

OpenStudy (anonymous):

\[8| x + \frac{ 3}{ 4 }| <2\]

OpenStudy (candy13106):

oh ok thats wat i thoight any way u add 3 over 4 to 8

OpenStudy (candy13106):

like this

OpenStudy (candy13106):